in a music award ceremony, 5 female singers are competing for the gold,silver, and bronze awards under "the best female singer".
1. if victoria gets the bronze, how many permutations are there for the winner in 3 awards?
2. if kate is one of the winner, how many permutations are there for the winner in the 3 awards?

Answer by ikleyn(52781)

.

( 1 )

Then there are 4 candidates for the gold


and 3 remaining candidates for silver.


In all,  4*3 = 12 permutations are possible.    ANSWER


( 2 )

If Kate won gold, then for the rest candidates there are 4*3 = 12 possible permutations.


If Kate won silver, then for the rest candidates there are 4*3 = 12 other possible permutations.


If Kate won bronze, then for the rest candidates there are 4*3 = 12 possible other permutations.


These three sets of permutations with 12 elements in each set, are DISGIUNT, 

so the final answer is  12+12+12 = 36 permutations.   


ANSWER.  In all, there are 36 possible permutations.